# Examining the Utility of Self-disclosure Types for Modeling Annotators of Social Norms

### v1

@@ F.7 C7: Sympathy & Support
The ’Sympathy & Support’ cluster contains comments which offer support and sympathy for the original poster. Almost all of the comments in this cluster contain "Not The Asshole" (NTA).
@@
Centeroid Comment: "Definitely NTA. I’m sorry you have to put up with that kind of behavior."
@@
Random Comment: "NTA. I’m just amazed that your sister is a 30 year old woman acting like that. "
F.8 C8: Family Disputes
@@
The ’Family Disputes’ cluster contains comments which discuss issues between family members. This usually occurs as arguments or somebody feeling slighted by another family member and asking Reddit for advice. This cluster is different from ’C4: Family Structure Change’ because it does not focus on a change in the structure of the family and rather on the issues and arguments between family members.
@@
Centeroid Comment: "ESH, apart from the sister. It’s understandable that you have these feelings, however, you’re portraying your sister as if she has the capacity to understand what she’s doing and does it to spite you, which makes you TA. Your parents sounds like they have neglected you, which is not ok, it was probably unintentional, but even so, it’s obviously had a deep impact on you. Your Dad’s comment was way out of line."
@@
Random Comment: "And your grandparents helped perpetuate this hurtful fraud. Fuck you really didn’t have anyone in your corner, it must have felt like you against the world. I am so sorry your family are such selfish hurtful fucks. NTA."
F.9 C9: Negative Affect
@@
The ’Negative Affect’ cluster contains comments in which people are disagreeing with something or talking about civility.
@@
Centeroid Comment: "I disagree with that. I like to judge the totality of a situation and not a mere instance. …"
Random Comment: "… I was objecting to the NAH I was replying to."
F.10 C10: Food & Meals
@@
The ’Food & Meals’ cluster contains comments which discuss topics related to food and eating.
@@
Centeroid Comment: "… You were my nightmare as a child. I was a terribly picky eater as a kid and going to sleepovers at friends’ houses was always a gamble because I was scared to try their food. Some people don’t like your food, that doesn’t make them an asshole, it just makes you not that great a cook. …"
@@
Random Comment: "YTA, I’m a chef and even I hate cooking dinner at home. I have two children so I cook when they are with me. It’s a lot harder to even cook for two people than it is for one. Double time preparing what ever ingredients are needed and probably more washing up at the end also. …"

### v2

@@ D.7 C7: Sympathy & Support
The Sympathy & Support cluster contains comments which offer support and sympathy for the original poster. Almost all of the comments in this cluster contain “Not The Asshole” (NTA).
@@
Centeroid Comment: “Definitely NTA. I’m sorry you have to put up with that kind of behavior.”
@@
Random Comment: “NTA. I’m just amazed that your sister is a 30 year old woman acting like that.”
D.8 C8: Family Disputes
@@
The Family Disputes cluster contains comments which discuss issues between family members. This usually occurs as arguments or somebody feeling slighted by another family member and asking Reddit for advice. This cluster is different from C4, Family Structure Change because it does not focus on a change in the structure of the family and rather on the issues and arguments between family members.
@@
Centeroid Comment: “ESH, apart from the sister. It’s understandable that you have these feelings, however, you’re portraying your sister as if she has the capacity to understand what she’s doing and does it to spite you, which makes you TA. Your parents sounds like they have neglected you, which is not ok, it was probably unintentional, but even so, it’s obviously had a deep impact on you. Your Dad’s comment was way out of line.”
@@
Random Comment: “And your grandparents helped perpetuate this hurtful fraud. Fuck you really didn’t have anyone in your corner, it must have felt like you against the world. I am so sorry your family are such selfish hurtful fucks. NTA.”
D.9 C9: Negative Affect
@@
The Negative Affect cluster contains comments in which people are disagreeing with something or talking about civility.
@@
Centeroid Comment: “I disagree with that. I like to judge the totality of a situation and not a mere instance. …”
Random Comment: “… I was objecting to the NAH I was replying to.”
D.10 C10: Food & Meals
@@
The Food & Meals cluster contains comments which discuss topics related to food and eating.
@@
Centeroid Comment: “… You were my nightmare as a child. I was a terribly picky eater as a kid and going to sleepovers at friends’ houses was always a gamble because I was scared to try their food. Some people don’t like your food, that doesn’t make them an asshole, it just makes you not that great a cook. …”
@@
Random Comment: “YTA, I’m a chef and even I hate cooking dinner at home. I have two children so I cook when they are with me. It’s a lot harder to even cook for two people than it is for one. Double time preparing what ever ingredients are needed and probably more washing up at the end also. …”